OpenStudy (anonymous):

Help Please! Reduce rational expression, 3x to the 2 power+14x-24 ______________________________ 18-9x-2x to the second power

OpenStudy (anonymous):

\[\frac{3 x^2+14 x-24}{-2 x^2-9 x+18}=\frac{(x+6) (3 x-4)}{(-x-6) (2 x-3)}=\frac{4-3 x}{2 x-3} \]